OpenStudy (jozelynw):

HELP!!!!!!! Algebra 1

OpenStudy (jozelynw):

\[(\sqrt{2}) (\sqrt[3]{2})\]

OpenStudy (jozelynw):

How I would I solve this equation

OpenStudy (calculusxy):

Do you mean: \[(\sqrt{2})(3\sqrt{2})\]

OpenStudy (jozelynw):

don't they mean the same thing

OpenStudy (jozelynw):

I guess so

OpenStudy (jhannybean):

no, they do not mean the same thing.

OpenStudy (jozelynw):

what is different between them?

OpenStudy (jhannybean):

\[\large \sqrt{2} \cdot \sqrt[3]{2} = 2^{1/2} \cdot 2^{1/3} = 2^{(1/2) + (1/3)}\]

OpenStudy (jhannybean):

Well, think about your exponent rule. \(\large \sqrt[n]{x^m} \implies x^{m/n}\)

OpenStudy (jozelynw):

I mean this one because, I started doing this but I don't know what to do next

OpenStudy (jhannybean):

Then when you have something like \(\sqrt{2} \cdot 3\sqrt{2}\) , you can easily factor out the common term, \(\sqrt{2}\) and add the multipliers. \[\sqrt{2} \cdot 3\sqrt{2} = \sqrt{2}(1+3) = 4\sqrt{2}\]

OpenStudy (jozelynw):

it's not that 1

OpenStudy (jhannybean):

Which one are you referring to, then?

OpenStudy (jozelynw):

The first 1

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @Jhannybean \[\large \sqrt{2} \cdot \sqrt[3]{2} = 2^{1/2} \cdot 2^{1/3} = 2^{(1/2) + (1/3)}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (jozelynw):

YES

OpenStudy (jhannybean):

And what is \(\frac{1}{2} +\frac{1}{3}\)?

OpenStudy (jozelynw):

2/5

OpenStudy (jozelynw):

it's not is it?

OpenStudy (jozelynw):

3/6 2/6

OpenStudy (jozelynw):

5/6 maybe

OpenStudy (jhannybean):

\[\frac{1}{2} + \frac{1}{3} = \frac{3}{6} + \frac{2}{6} = \frac{ 2+3}{6} = \frac{5}{6}\]

OpenStudy (nincompoop):

calling jhan right now

OpenStudy (jozelynw):

oh i was right ok

OpenStudy (nincompoop):

wash the dishes, jhan they are piling up

OpenStudy (jozelynw):

ok so whats next?

OpenStudy (jhannybean):

Therefore you can write the fraction exponent as: \(2^{5/6} \iff \sqrt[6]{2^5}\)

OpenStudy (jozelynw):

Ok thank you I get it.